# ENTROPY-REGULARIZED GRADIENT ESTIMATORS FOR APPROXIMATE BAYESIAN INFERENCE

### v1

@@ 4.1 Bivariate Gaussian Mixture
For the case shown, methods like SVGD, kde-WGD, sse-WGD and SGD do not find all the modes. The last image in the third row shows -ERGD is able to explore enough to find all the modes. It is to be kept in mind that -ERGD can fail to converge. ERGD using a linear schedule still fails to find all the modes. D’Angelo and Fortuin [2021a] suggested that a linear schedule can lead to slower training dynamics and proposed a cyclic tanh schedule. ERGD with a tanh schedule can find some of the modes. Since -ERGD can explore better, the second last figure in the last row uses -ERGD to find an initial distribution which is then converged to the final solution using a linear schedule. So instead of using a different schedule, -ERGD can be used to find such an initial distribution while keeping the convergence properties of ERGD.
@@ B.2 Model-Based RL
Training Epochs | 200 | Number of training epochs |

### v2

@@ 4.1 Bivariate Gaussian Mixture
It is important to note that s-ERGD may fail to converge in certain cases. Using ERGD with a linear schedule still fails to capture all modes of the distribution. D’Angelo and Fortuin [2021a] suggests that a linear schedule could result in slower training dynamics and proposed the use of a cyclic tanh schedule as an alternative. While ERGD with a tanh schedule can capture some of the modes, it does not necessarily guarantee complete mode discovery. On the other hand, s-ERGD is capable of exploring the distribution more effectively. Therefore, in the second-to-last figure of the final row, s-ERGD is employed to obtain an initial distribution, which is subsequently converges to the final solution using a linear schedule. This approach allows s-ERGD to serve as a way to have an initial distribution, while having the convergence properties associated with ERGD.
@@ B.2 Model-Based RL
| Training Epochs | 100 | Number of training epochs |